# Quantum simulating continuum field theories with large-spin lattice models

### v1

@@ VI.1 Choice and interpretation of the perturbation
with $p$ the strength of the perturbation. The corresponding lattice version of the perturbed Hamiltonian $\hat{H}_{\text{latt,p}}$ could be realized, e.g., in a Rydberg setup using both a microwave coupling with $\kappa=1$ and a ponderomotive coupling with $\kappa=2$ [see App. A]. The microscopic parameter $\lambda^{\prime}_{2}$ is then fixed according to Eq. (9), while $\lambda^{\prime}_{1}=p\cdot\lambda_{2}^{\prime}$.
@@
From this perspective, we view the double-frequency sine-Gordon model Eq. (34) as a rough approximation of the massive sG model [18] with a potential $V(\hat{\varphi})\leavevmode\nobreak\ \sim\leavevmode\nobreak\ +\frac{M_{0}^{2}}{2}\hat{\varphi}^{2}\leavevmode\nobreak\ -\leavevmode\nobreak\ p\frac{M_{0}^{2}}{\beta^{2}}\cos(\frac{\beta\hat{\varphi}}{2})$. As mentioned in Sec. II.3, the latter is known to be the dual bosonized version of the massive Schwinger model, i.e., quantum electrodynamics in one spatial dimension. Motivated by this relation, we will interpret the scattering dynamics qualitatively in terms of “electric fields” and “electron-positron” pair production (= excitations with negative/positive charge).
Figure 7: Beyond integrability. (a) Interaction potential $V(\hat{\varphi})\beta^{2}/M_{0}^{2}$ of Eq. (35) for $p=0$ (gray line) and $p=0.5$ (green line). (b) Energy density $\varepsilon_{i}$ (in sine-Gordon units) of the soliton-antisoliton static pair (dashed black) for the unperturbed $p=0$ (solid gray) and the perturbed $p=0.5$ (solid green). The dashed black line represents the classical soliton-antisoliton profile, while black dots capture the quasi-particle position in the unperturbed case. The simulation parameters are $N=101,J=22,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2,\Delta_{\text{s}\bar{\text{s}}}/a=40$. (c) resp. (e) Topological charge density $\rho$ with quasi-particles trajectories (solid black) as a function of time $t$ and site $i$ for $p=0.5$ resp. $p=1.5$. (d) Von Neumann entanglement entropy, in units of $\log(2)$, with quasi-particles trajectories for $p=0.5$. (f) Expectation value of $\sin(\beta\hat{\varphi})$ as a function of time $t$ and site $i$ for $p=1.5$. This observable can be interpreted as an “electric” field [see main text].
@@ VI.2 Numerical results
Figure 8: Dynamics of a soliton-antisoliton pair in the perturbed sine-Gordon model. We investigate the dynamics of an initially static soliton and antisoliton for the parameters $J=22,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2,\Delta_{\text{s}\bar{\text{s}}}/a=40$. The perturbation strength $p$ of the perturbed sine-Gordon model is increasing from left to right: $p=0.5,1.0,1.5,2.5,3.5$. (First row) Topological charge density $\rho$: the red color represents a positive density (soliton), the blue a negative one (antisoliton). (Second row) Expectation value of the operator $\sin(\beta\hat{\varphi})$, as a proxy for the “electric” field. (Third row) Normalized energy density, obtained by subtracting the ground-state energy density under the perturbed Hamiltonian. For each perturbation strength the color map goes from the respective minimum to the maximal value. (Fourth row) Von Neumann entanglement entropy $S_{\text{vN}}$ for each bond along the chain.
@@
This picture radically changes upon increasing the perturbation strength as shown in Fig. 8. At $p\sim 1.5$ the periodic oscillations and collisions disappear. Instead, we observe the production of new particles in the middle of the chain, as evidenced by the topological charge density $\rho$ in Fig. 7(e). Moreover, after the scattering event the observable $\langle\sin(\beta\hat{\varphi})\rangle$ displays pronounced oscillations in the middle of the chain [see Fig. 7(f)]. To put this into context, we note that in the Schwinger model strong electric fields lead to repeated particle pair-production events inducing plasma oscillations of the electric field [73]. In the dual massive sine-Gordon model, it is precisely the observable $\sin(\beta\hat{\varphi})\approx\beta\hat{\varphi}$ that plays the role of the electric field. The observed behaviour is therefore qualitatively similar to a breaking of the initial electric flux string and the expected subsequent plasma oscillations. Moreover, the half-chain von Neumann entanglement entropy at $p=1.5$ displays recurring local maxima, which we may interpret as a signature of the creation of new particle pairs. Note that for all observables the late-time dynamics is clearly affected by the reflection of excitations at the boundaries of the finite system.
@@ VII Conclusions and outlook
Our results form the basis for studying non-equilibrium real-time dynamics of QFTs with large-spin models. Relevant examples for future work include the decay of the false vacuum [8, 51, 52, 53, 43], or a more extensive study of scattering in non-integrable models [75, 76, 77, 78, 79]. While such studies are possible in one spatial dimension using classical tensor-network techniques, an experimental implementation of our approach in higher spatial dimensions has the potential to reach uncharted territory.
From an experimental perspective, the required spin lengths are not too large ($J\lesssim 20$), making it interesting to further investigate the feasibility of implementing our proposal with Rydberg atom arrays [18, 23]. One important point here is the (ir)relevance of long-range tails of the dipole-dipole interactions for the continuum physics [80]. A possible verification of an experimental implementation of the desired QFT could be achieved using Hamiltonian learning techniques, which were recently adapted to the field theory setting [81].
While our approach was motivated by analog quantum simulations, our results are also of interest for digital quantum simulations based on qudits [25, 16], as well as hybrid digital-analog approaches [82, 24]. In the context of qubit vs. qudit computers, and given the limited resources of existing hardware, an important outstanding question is how to make optimal use of available hardware. In the case of QFTs, this question is especially intriguing because we are only interested in universal continuum physics independent of irrelevant microscopic details. As the numerical results in this work have been achieved with remarkably low bond dimensions, comparable with analogous spin-$1/2$ models, we speculate that the large-spin approach provides some level of encoding, which could imply a natural robustness against errors. This conjecture could be tested in a detailed resource comparison of, e.g., a spin-$1/2$ XXZ realization of the sine-Gordon model vs. the approach presented here, including the effects of experimental imperfections, such as finite coherence or gate fidelities, which we leave for future work.
@@ Acknowledgements
Appendix A Description of large-spin Rydberg atom arrays
@@
Appendix B Mapping the lattice Hamiltonian to the continuum scalar QFT
@@
allows the simulation of the continuum scalar QFTs Hamiltonian $\hat{H}_{\text{QFT}}$ in Eq. (8) upon proper choice of the parameters. We begin by considering a single $\kappa$ and choose $\Omega=\Delta=0$. In order to employ the identification between spin and QFT operators Eq. (3), we further rescale the lattice parameters as $V_{\text{nn}}^{\prime}=V_{\text{nn}}J(J+1)$ and $\lambda_{\kappa}^{\prime}=\lambda_{\kappa}[J(J+1)]^{\kappa/2}$. Upon rescaling, the Ising term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$ acquires an additional factor $1/J(J+1)$ and becomes negligible in the large spin-length limit $J\rightarrow\infty$ [see App. E.1]. We thus obtain the discrete lattice version of the scalar QFTs Hamiltonian
@@
where we sum over the spatial derivatives $[\partial_{x_{k}}\hat{\varphi}(x)]^{2}$ in each of the $d$ directions. The implementation of the continuum QFT from the Rydberg lattice Hamiltonian is only valid in the limit of large spin length $J\rightarrow\infty$, large atom number $N\rightarrow\infty$, as well as in the continuum limit $M_{0}^{\prime}\rightarrow 0$ [see discussion in the main text].
Appendix C Estimation of the continuum limit of observables
@@
As discussed in the main text, the analog quantum simulation of scalar QFTs on a lattice is only valid in the continuum limit and thus, we need to extrapolate the value and uncertainty of relevant observables in the limit $J,N\rightarrow\infty$. To this end, we run numerical simulations for fixed QFT parameters $\beta^{2},M_{0}^{\prime}$ at several $J$ and $N$, thus obtaining a set of data points $\{O_{i}\}$. We then fix $N$ and consider a minimal subset of data points, corresponding to the numerical results at the largest $J$ values simulated. We fit this minimal subset with a function $h(x)$ linear in $x=1/J(J+1)$, thus obtaining a first estimate $O_{\alpha}$ for the asymptotic limit $J\rightarrow\infty$. We repeat this procedure for increasingly larger subsets of $\{O_{i}\}$ by adding data points at smaller $J$ values, and thus obtain new predictions for the asymptotic value $O_{\alpha}$ ($\alpha=1,\dots,N_{\text{fits}}$, with $N_{\text{fits}}$ the number of subsets considered). If we do not have an estimate for the error $\Delta O_{i}$ of the data point $O_{i}$, we choose as our best guess $\bar{O}_{\alpha}$ for the asymptotic limit the median value of the set $\{O_{\alpha}\}$, and we estimate the uncertainty $\Delta O_{\alpha}$ as half of the difference between the largest and smallest values $\{O_{\alpha}\}$. Instead, if we have already an estimate for the error $\Delta O_{i}$, the best guess $\bar{O}_{\alpha}$ is determined according to a cumulative distribution function $f_{\alpha}$, introduced in Ref. [83],
@@
Appendix D Conformal field theory and next-nearest-neighbor interaction
@@
In Fig. 9(a) we compare the continuum results for both the energy gap $\Delta E_{\text{latt}}$ (in lattice units) and the central charge $c$ in two cases: for nearest-neighbors only, and also the next-nearest-neighbor term of the dipole-dipole interaction. We observe that higher orders in the dipole-dipole interaction only modify the value of the mass gap $\Delta E_{\text{latt}}$ in the gapped region (where the lattice model does not correctly simulate the CFT), while the central charge is only minimally modified in the gapless phase. We thus expect that the long-range interaction does not extend the simulability region of the model, yet it modifies the value of the Luttinger parameter $K$. Note that the central charge $c$ is extrapolated upon fitting with Eq. (17) the von Neumann entanglement entropy $S_{\text{vN}}$ at different cut positions along the chain [see Fig. 9(b)]. The continuum result is then determined from a linear fit in $1/J(J+1)$ and a subsequent fit in $1/N$, as described in App. C.
@@
We extrapolate the value of the Luttinger parameter $K$ from the connected vertex-vertex correlator $\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(i+r)}\rangle_{\text{c}}/J(J+1)$ by linearly fitting the logarithm of the correlator against the logarithm of the separation $r$ between the two sites considered according to Eq. (III.2.1). We choose the fitting interval in such a way to avoid boundary effects (we neglect the 15 sites closest to both ends), and to fit over linear regions only [see inset in Fig. 9(c)]. The Luttinger parameter $K$ displays a tiny dependence on the site $i$ at which it is calculated due to finite-size effects; therefore, we consider as our best guess the half-chain result $i=N/2$. In Fig. 9(d) we compare the numerical results for the Luttinger parameter $K$ with the theoretical prediction Eq. (19), and notice that it displays some renormalization w.r.t. the microscopic mapping Eq. (19). We expect the renormalization to emerge due to the truncation of the cosine term arising from the dipole-dipole interaction. As expected [see App. E.2], the renormalization increases for decreasing $V_{\text{nn}}^{\prime}$, but it is partially compensated by the NNN interaction term. Indeed, the value $K$ obtained upon considering also next-nearest-neighbors terms is very close to the microscopic prediction Eq. (19) in the simulated regime.
Appendix E Equilibrium properties of the sine-Gordon model
@@
In this Appendix we discuss additional results regarding the equilibrium properties of the sine-Gordon model. In App. E.1 we discuss the impact of neglected next-nearest-neighbors dipole-dipole interaction and Ising terms on the simulated sG model. In App. E.2 we discuss the role of the neglected higher-order terms in the cosine expansion and how they impact the renormalization of the parameter $\beta^{2}$. In App. E.3 we illustrate two methods to determine the renormalization of the mass parameter $M_{0}^{\prime}$. In App. E.4 we discuss the Lukyanov-Zamolodchikov conjecture.
E.1 Neglected terms in the dipole-dipole interaction
@@
The sG model can be simulated with the 1D lattice Hamiltonian $\hat{H}_{\text{latt}}$ [Eq. (II.1)], obtained from the Rydberg Hamiltonian $\hat{H}_{\text{Ryd}}$ [Eq. (36)] by neglecting higher-order dipole-dipole interaction term $V_{ij}$ ($j>i+1$) and the Ising contribution $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$. The latter approximation is justified in the asymptotic limit $J\rightarrow\infty$: using the identification Eq. (3) between lattice spin operators and the continuum QFT operators, as well as by introducing the parameter $V_{ij}^{\prime}=V_{ij}J(J+1)$, the term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$ is rescaled by the factor $1/J(J+1)$, which makes the Ising term vanish in the large-spin limit $J\rightarrow\infty$ [see App. B].
@@
In Fig. 10(e) we compare the continuum results of both the mass gap and the vertex operator at $V_{\text{nn}}^{\prime}/\chi=100/\pi^{2}$ and for different $\sqrt{\lambda_{\kappa}^{\prime}/\chi}$ (i.e., $M_{0}^{\prime}$). For the gap the data points of the three cases agree with each other, while for the vertex operator we observe some deviations between NN and NNN, which increase for decreasing $M_{0}^{\prime}$. This deviation leads to different renormalized $\beta^{2}$ values [see Fig. 10(f)], which, as discussed in the main text, we extrapolate from the log-log fit of mass gap vs. vertex operator. We notice that the NNN interaction reduces the renormalization [see Fig. 10(f)], in agreement with the CFT case [see App. D]. This behaviour indicates a counteracting effect of the long-range dipole-dipole interaction on the neglected terms in the expansion of the cosine [see App. E.2 for further details].
@@
Figure 10: Effect of neglected interactions on the sG model simulation. Comparison of numerical results obtained from DMRG calculations of the lattice model $\hat{H}_{\text{latt}}$ upon considering only nearest-neighbor dipole-dipole interaction (NN, blue dots), by adding the Ising term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(i+1)}$ (NN + Ising, green diamonds), or by adding next-nearest-neighbor interactions (NNN, orange squares). We choose $\kappa=1$ and unless varying over the corresponding parameters, we consider $N=100,\lambda_{\kappa}^{\prime}=1/\pi^{2},\chi=1$ and $V_{\text{nn}}^{\prime}=400/\pi^{2}$ in (a)-(b) resp. $V_{\text{nn}}^{\prime}=100/\pi^{2}$ in (c)-(e). (a), (c) Energy gap in lattice units $\Delta E_{\text{latt}}$ and corresponding fit in $1/J(J+1)$ for the different interaction terms. (b), (d) The asymptotic value extracted in the large-spin length limit $J\rightarrow\infty$ is plotted for different system sizes $N$ and fitted against $1/N$. (e) Continuum results for the lattice gap $\Delta E_{\text{latt}}$ and (inset) the vertex operator $\langle e^{i\beta\hat{\varphi}}\rangle=\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)}$ as a function of the microscopic parameter ratio $\sqrt{\lambda_{\kappa}^{\prime}/\chi}$. (f) Renormalized $\beta^{2}$ value against the microscopic ratio $\sqrt{\chi/V_{\text{nn}}^{\prime}}$. The dashed black line represents the microscopic prediction according to Eq. (9).
E.2 $\beta^{2}$-renormalization and $\kappa$-dependence

### v2

@@ VI.1 Choice and interpretation of the perturbation
with $p$ the strength of the perturbation. The corresponding lattice version of the perturbed Hamiltonian $\hat{H}_{\text{latt,p}}$ could be realized, e.g., in a Rydberg setup using both a microwave coupling with $\kappa=1$ and a ponderomotive coupling with $\kappa=2$ [see App. B]. The microscopic parameter $\lambda^{\prime}_{2}$ is then fixed according to Eq. (9), while $\lambda^{\prime}_{1}=p\cdot\lambda_{2}^{\prime}$.
@@
[…]
@@ VI.2 Numerical results
We initialize our system by imprinting the desired phase profile, this time on the ground state of the quantum sine-Gordon Hamiltonian realized with $\kappa=2$. To deal with the new $4\pi$-symmetry, we first check that the spin is pointing in +x+x-direction (if instead it is pointing in x-x-direction, we rotate it accordingly), before applying the site-dependent $z$-rotations with angle $\beta\varphi_{\text{s}\bar{\text{s}}}^{(i)}/2$, such that the expectation value of the phase profile $\langle\beta\hat{\varphi}\rangle=\arg\langle(\hat{J}_{+}/\sqrt{J(J+1)})^{2}\rangle$ coincides with the desired classical solution. We then evolve the prepared semi-classical soliton-antisoliton superposition under the perturbed Hamiltonian $\hat{H}_{\text{sG,p}}$ for several perturbation strengths $0.5\leavevmode\nobreak\ \leq\leavevmode\nobreak\ p\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 3.5$.
@@
[…]
@@ VII Conclusions and outlook
Our results form the basis for studying non-equilibrium real-time dynamics of QFTs in the continuum limit with large-spin models. Relevant examples for future work include the decay of the false vacuum [9, 67, 68, 69, 59], or a more extensive study of scattering in non-integrable models [91, 37, 92, 93, 94, 95]. While such studies are possible in one spatial dimension using classical tensor-network techniques, an experimental implementation of our approach in higher spatial dimensions has the potential to reach uncharted territory.
From an experimental perspective, the required spin lengths are not too large ($J\lesssim 20$), making it interesting to further investigate the feasibility of implementing our proposal with Rydberg atom arrays [33, 41]. Such a setup could also make use of quantum state verification protocols [96] to perform a self-consistent convergence test with respect to the spin length within a single experiment. Another important point in this atomic setup is the (ir)relevance of long-range tails of the dipole-dipole interactions for the continuum physics [97].
While our approach was motivated by analog quantum simulations, our results are also of interest for digital quantum simulations based on qudits [42, 29], as well as hybrid digital-analog approaches [98, 32]. In this context, an important outstanding question is the resource trade-off between the local Hilbert space dimension (spin length or qudit size) and the total system size (i.e. number of spins or qudits). In the case of QFTs, this question is especially intriguing because we are only interested in universal continuum physics independent of irrelevant microscopic details. As the numerical results in this work have been achieved with remarkably low bond dimensions, comparable with analogous spin-$1/2$ models, we speculate that the large-spin approach provides some level of encoding, which could imply a natural robustness against errors. This conjecture could be tested in a detailed resource comparison of, e.g., a spin-$1/2$ XXZ realization of the sine-Gordon model vs. the approach presented here, including the effects of experimental imperfections, such as finite coherence or gate fidelities, which we leave for future work.
Throughout this work we demonstrated quantitative agreement of our approach by comparing the extrapolated numerical results to exact analytical QFT prediction. In the case of an unknown QFT to be simulated, it will be essential to perform an error-bounding analysis, especially in the most interesting regime where classical simulations are infeasible [5]. While for continuum QFTs a naive analysis based on operator norms is not possible due the unboundedness of the involved operators, we note that this issue could be avoided by projecting to finite-energy subspaces [99, 22]. A possible verification of an analog experimental implementation of the desired QFT could be achieved using Hamiltonian learning techniques, which were recently adapted to the field theory setting [100].
@@ Acknowledgements
Appendix A Overview table
In order to quickly locate plots in the main text, below we provide a table of the investigated observables and their corresponding figures for the three studied models.
Table 1: Investigated observables for each QFT model and corresponding figure.
  QFT Model   Observable   Figure
   Luttinger liquid CFT   Central charge $c$   Inset 3a
  Energy gap $\Delta E_{\mathrm{latt}}$   3a
  $\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(i\pm r)}\rangle/\sqrt{J(J+1)}$   Inset 3b
  Luttinger parameter $K$   3b
  sine-Gordon model   Renormalized $\beta^{2}$   4a
  Ground-state energy density $E_{0,\mathrm{latt}}$   4b
  Energy gap $\Delta E_{\mathrm{sG}}$   Inset 4b
  Vertex operator $\langle e^{i\beta\hat{\varphi}}\rangle$   4c
  Higher powers of the vertex operator $\langle e^{in\beta\hat{\varphi}}\rangle$   4d
  Topological charge density $\rho$   5c
  $\langle\cos{\beta\hat{\varphi}}\rangle$   5d
  von Neumann entanglement entropy $S_{\mathrm{vN}}$   Inset 5e
  Excess lattice energy density $\bar{\varepsilon}$   6a
  Position shift $\delta x$   6b
   Perturbed sine-Gordon model   Topological charge density $\rho$   7c / 7e / 8 (1st row)
  von Neumann entanglement entropy $S_{\mathrm{vN}}$   7d / 8 (4th row)
  $\langle\sin{\beta\hat{\varphi}}\rangle$   7f / 8 (2nd row)
  Normalized energy density   8 (3rd row)
Appendix B Description of large-spin Rydberg atom arrays
@@
Appendix C Mapping the lattice Hamiltonian to the continuum scalar QFT
@@
allows the simulation of the continuum scalar QFTs Hamiltonian $\hat{H}_{\text{QFT}}$ in Eq. (8) upon proper choice of the parameters. We begin by considering a single $\kappa$ and choose $\Omega=\Delta=0$. In order to employ the identification between spin and QFT operators Eq. (3), we further rescale the lattice parameters as $V_{\text{nn}}^{\prime}=V_{\text{nn}}J(J+1)$ and $\lambda_{\kappa}^{\prime}=\lambda_{\kappa}[J(J+1)]^{\kappa/2}$. Upon rescaling, the Ising term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$ acquires an additional factor $1/J(J+1)$ and becomes negligible in the large spin-length limit $J\rightarrow\infty$ [see App. F.1]. We thus obtain the discrete lattice version of the scalar QFTs Hamiltonian
@@
where we sum over the spatial derivatives $\partial_{x_{k}}\hat{\varphi}(x)]^{2}$ in each of the $d$ directions. The implementation of the continuum QFT from the Rydberg lattice Hamiltonian is only valid in the limit of large spin length $J\rightarrow\infty$, large atom number $N\rightarrow\infty$, as well as in the continuum limit $M_{0}^{\prime}\rightarrow 0$ [see discussion in the main text].
Appendix D Estimation of the continuum limit of observables
@@
As discussed in the main text, the analog quantum simulation of scalar QFTs on a lattice is only valid in the continuum limit and thus, we need to extrapolate the value and uncertainty of relevant observables in the limit $J,N\rightarrow\infty$. To this end, we run numerical simulations for fixed QFT parameters $\beta^{2},M_{0}^{\prime}$ at several $J$ and $N$, thus obtaining a set of data points $\{O_{i}\}$. We then fix NN and consider a minimal subset of data points, corresponding to the numerical results at the largest $J$ values simulated. We fit this minimal subset with a function $h(x)$ linear in $x=1/J(J+1)$, thus obtaining a first estimate $O_{\alpha}$ for the asymptotic limit $J\rightarrow\infty$. We repeat this procedure for increasingly larger subsets of $\{O_{i}\}$ by adding data points at smaller $J$ values, and thus obtain new predictions for the asymptotic value $O_{\alpha}$ ($\alpha=1,\dots,N_{\text{fits}}$, with $N_{\text{fits}}$ the number of subsets considered). If we do not have an estimate for the error $\Delta O_{i}$ of the data point $O_{i}$, we choose as our best guess $\bar{O}_{\alpha}$ for the asymptotic limit the median value of the set $\{O_{\alpha}\}$, and we estimate the uncertainty $\Delta O_{\alpha}$ as half of the difference between the largest and smallest values $\{O_{\alpha}\}$. Instead, if we have already an estimate for the error $\Delta O_{i}$, the best guess $\bar{O}_{\alpha}$ is determined according to a cumulative distribution function $f_{\alpha}$, introduced in Ref. [101],
@@
Appendix E Conformal field theory and next-nearest-neighbor interaction
@@
In Fig. 9(a) we compare the continuum results for both the energy gap $\Delta E_{\text{latt}}$ (in lattice units) and the central charge $c$ in two cases: for nearest-neighbors only, and also the next-nearest-neighbor term of the dipole-dipole interaction. We observe that higher orders in the dipole-dipole interaction only modify the value of the mass gap $\Delta E_{\text{latt}}$ in the gapped region (where the lattice model does not correctly simulate the CFT), while the central charge is only minimally modified in the gapless phase. We thus expect that the long-range interaction does not extend the simulability region of the model, yet it modifies the value of the Luttinger parameter $K$. Note that the central charge $c$ is extrapolated upon fitting with Eq. (17) the von Neumann entanglement entropy $S_{\text{vN}}$ at different cut positions along the chain [see Fig. 9(b)]. The continuum result is then determined from a linear fit in $1/J(J+1)$ and a subsequent fit in $1/N$, as described in App. D.
@@
We extrapolate the value of the Luttinger parameter $K$ from the connected vertex-vertex correlator $\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(i+r)}\rangle_{\text{c}}/J(J+1)$ by linearly fitting the logarithm of the correlator against the logarithm of the separation $r$ between the two sites considered according to Eq. (III.2.1). We choose the fitting interval in such a way to avoid boundary effects (we neglect the 15 sites closest to both ends), and to fit over linear regions only [see inset in Fig. 9(c)]. The Luttinger parameter $K$ displays a tiny dependence on the site $i$ at which it is calculated due to finite-size effects; therefore, we consider as our best guess the half-chain result $i=N/2$. In Fig. 9(d) we compare the numerical results for the Luttinger parameter $K$ with the theoretical prediction Eq. (19), and notice that it displays some renormalization w.r.t. the microscopic mapping Eq. (19). We expect the renormalization to emerge due to the truncation of the cosine term arising from the dipole-dipole interaction. As expected [see App. F.2], the renormalization increases for decreasing $V_{\text{nn}}^{\prime}$, but it is partially compensated by the NNN interaction term. Indeed, the value $K$ obtained upon considering also next-nearest-neighbors terms is very close to the microscopic prediction Eq. (19) in the simulated regime.
Appendix F Equilibrium properties of the sine-Gordon model
@@
In this Appendix we discuss additional results regarding the equilibrium properties of the sine-Gordon model. In App. F.1 we discuss the impact of neglected next-nearest-neighbors dipole-dipole interaction and Ising terms on the simulated sG model. In App. F.2 we discuss the role of the neglected higher-order terms in the cosine expansion and how they impact the renormalization of the parameter $\beta^{2}$. In App. F.3 we illustrate two methods to determine the renormalization of the mass parameter $M_{0}^{\prime}$. In App. F.4 we discuss the Lukyanov-Zamolodchikov conjecture.
F.1 Neglected terms in the dipole-dipole interaction
@@
The sG model can be simulated with the 1D lattice Hamiltonian $\hat{H}_{\text{latt}}$ [Eq. (II.1)], obtained from the Rydberg Hamiltonian $\hat{H}_{\text{Ryd}}$ [Eq. (36)] by neglecting higher-order dipole-dipole interaction term $V_{ij}$ ($j>i+1$) and the Ising contribution $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$. The latter approximation is justified in the asymptotic limit $J\rightarrow\infty$: using the identification Eq. (3) between lattice spin operators and the continuum QFT operators, as well as by introducing the parameter $V_{ij}^{\prime}=V_{ij}J(J+1)$, the term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}$ is rescaled by the factor $1/J(J+1)$, which makes the Ising term vanish in the large-spin limit $J\rightarrow\infty$ [see App. C].
@@
In Fig. 10(e) we compare the continuum results of both the mass gap and the vertex operator at $V_{\text{nn}}^{\prime}/\chi=100/\pi^{2}$ and for different $\sqrt{\lambda_{\kappa}^{\prime}/\chi}$ (i.e., $M_{0}^{\prime}$). For the gap the data points of the three cases agree with each other, while for the vertex operator we observe some deviations between NN and NNN, which increase for decreasing $M_{0}^{\prime}$. This deviation leads to different renormalized $\beta^{2}$ values [see Fig. 10(f)], which, as discussed in the main text, we extrapolate from the log-log fit of mass gap vs. vertex operator. We notice that the NNN interaction reduces the renormalization [see Fig. 10(f)], in agreement with the CFT case [see App. E]. This behaviour indicates a counteracting effect of the long-range dipole-dipole interaction on the neglected terms in the expansion of the cosine [see App. F.2 for further details].
@@
Figure 10: Effect of neglected interactions on the sG model simulation. Comparison of numerical results obtained from DMRG calculations of the lattice model $\hat{H}_{\text{latt}}$ upon considering only nearest-neighbor dipole-dipole interaction (NN, blue dots), by adding the Ising term $\hat{J}_{z}^{(i)}\hat{J}_{z}^{(i+1)}$ (NN + Ising, green diamonds), or by adding next-nearest-neighbor interactions (NNN, orange squares). We choose $\kappa=1$ and unless varying over the corresponding parameters, we consider $N=100,\lambda_{\kappa}^{\prime}=1/\pi^{2},\chi=1$ and $V_{\text{nn}}^{\prime}=400/\pi^{2}$ in (a)-(b) resp. $V_{\text{nn}}^{\prime}=100/\pi^{2}$ in (c)-(e). (a), (c) Energy gap in lattice units $\Delta E_{\text{latt}}$ and corresponding fit in $1/J(J+1)$ for the different interaction terms. (b), (d) The asymptotic value extracted in the large-spin length limit $J\rightarrow\infty$ is plotted for different system sizes NN and fitted against $1/N$. (e) Continuum results for the lattice gap $\Delta E_{\text{latt}}$ and (inset) the vertex operator $\langle e^{i\beta\hat{\varphi}}\rangle=\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)}$ as a function of the microscopic parameter ratio $\sqrt{\lambda_{\kappa}^{\prime}/\chi}$. (f) Renormalized $\beta^{2}$ value against the microscopic ratio $\sqrt{\chi/V_{\text{nn}}^{\prime}}$. The dashed black line represents the microscopic prediction according to Eq. (9).
F.2 $\beta^{2}$-renormalization and $\kappa$-dependence